Electricity costs $0.10 per kilowatt-hour. How does it cost a family each year if they use 250 kilowatt-hours per month?


1
Expert's answer
2021-01-14T10:40:26-0500

Let's calculate how the electricity costs a family each year:


"Cost = Rate\cdot E \cdot \dfrac{12\ month}{1\ year},""Cost = 0.10\ \dfrac{\$}{kWh}\cdot 250\ \dfrac{kWh}{month} \cdot \dfrac{12\ month}{1\ year}=300\ \dfrac{\$}{year}."

Answer:

"Cost = 300\ \dfrac{\$}{year}."